In a resonance column experiment, a tuning fork of frequency 400 Hz is used. The first resonance is observed when the air column has a length of 20.0 cm and the second resonance is observed when the air column has a length of 62.0 cm. (a) Find the speed of sound in air. (b) How much distance above the open end does the pressure node form?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Length of the wire = 1.0 m For fundamental frequency λ/2 = l ⇒ λ = 2l = 2 * 1 = 2 m Here given n = 3.8 km/s = 3800 m/s We know ⇒ v = nλ⇒ n = 3800 / 2 = 1.9 kH. So standing frequency between 20 Hz and 20 kHz which will be heard are = n × 1.9 kHz where n = 0, 1, 2, 3, … 10.